SUBSTITUTION STRUCTURES OF MULTIPLE SILICON-CONTAINING SPECIES BY CHIRPED PULSE FTMW SPECTROSCOPY

\underline{NATHAN A. SEIFERT}, SIMON LOBSIGER, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904-4319; GAMIL A. GUIRGIS, JASON S. OVERBY, Department of Chemistry & Biochemistry, College of Charleston, Charleston, SC 29424 USA; JAMES R. DURIG, Department of Chemistry, University of Missouri-Kansas City, Kansas City, MO 64110 USA.

The structures of CH_{3}SiHF-NCO, 1-X-silacyclopropane (X = cyano, isocyanato), 1,1,3,3-tetrafluoro-1,3-disilacyclopentane and its hydrogen analogue (1,3-disilacyclopentane), and 1-isocyanato-silacyclohexane have been studied by chirped pulse FTMW spectroscopy in the 6-18 GHz band. Multiple conformers for some of the species were also detected: anti and gauche for both silacyclopropyl species, and axial and equatorial for the silacyclohexane. Heavy atom substitution structures were determined, with all possible single ^{13}C, ^{29}Si/^{30}Si and most ^{15}N isotopologues assigned in natural abundance. Nitrogen hyperfine and distortion parameters for all species have been determined, and the barrier for methyl internal rotation for CH_{3}SiHF-NCO has been determined as 481(20) cm^{−1}, close to the B3LYP/6-311++g(d,p) barrier of 450 cm^{−1}. A summary of the microwave and structural results for the aforementioned molecules will be presented. In addition, emphasis will be placed on the use of previously discussed^{a}^{b} automated fitting techniques as a means of efficient and fast assignment of isotopologues in spectra with increasingly large line densities.

aA. L. Steber, J. L. Neill, M. T. Muckle, B. H. Pate, D. F. Plusquellic, V. Lattanzi, S. Spezzano, M. C. McCarthy. 65th OSU Int. Symp. On Mol. Spectrosc., Columbus, OH, 2010, TC10.

bE. B. Kent, M. N. McCabe, M. A. Phillips, B. P. Gordon, S. T. Shipman. 66th OSU Int. Symp. On Mol. Spectrosc., Columbus, OH, 2011, RH01.